OpenStudy (anonymous):

If 2 +squarerootof 3 is a polynomial root, name another root of the polynomial and explain how you know it must also be a root

OpenStudy (anonymous):

no

Parth (parthkohli):

Seen the quadratic formula?

OpenStudy (anonymous):

its -2-root3.. See the polynomial is going to be 2+root3 multiplied by itself and the product of it is going to have two square roots of which -2-root3 is one of them

Parth (parthkohli):

Look how it says that if \(s\) and \(r\) are roots, then\[s = \frac{-b + \sqrt{b^2 - 4ac}}{2a}\]and\[r = \frac{-b - \sqrt{b^2 - 4ac}}{2a}\]

Parth (parthkohli):

So what would happen if we take \(s = 2 + \sqrt{3}\)?

OpenStudy (anonymous):

@ Parth it says polynomial not quadtraic